OpenStudy (anonymous):

A block of mass m slides down on rough inclined plane of angle x with the horizontal with an acceleration "a" along the plane the value of reaction is?

OpenStudy (anonymous):

@experimentX @ghazi @dumbcow @mukushla

OpenStudy (ghazi):

well you have to consider reaction of that mass at an instant ...so it will be = \[m*g*\cos \theta\] where theta is the inclination ....

OpenStudy (ghazi):

okay wait

OpenStudy (ghazi):

didn't see acceleration

OpenStudy (anonymous):

The answer should be \[m \sqrt{g^2 + ^2 - 2 a g sinx}\]

OpenStudy (ghazi):

i think your answer needs modification

OpenStudy (anonymous):

\[m g cosx\]

OpenStudy (anonymous):

i got this...

OpenStudy (anonymous):

Wat can be Done here,,,lol

OpenStudy (ghazi):

well mgcosx is the answer when block is stationary ..it's moving so you've to resolve resultant acceleration first

OpenStudy (anonymous):

oh......)) Can u show..that

OpenStudy (ghazi):

yea wait...

OpenStudy (anonymous):

Ok....

OpenStudy (ghazi):

 find R first by parallelogram law of vector addition and then use cos component of that R to multiply with m*g to find the reaction force

OpenStudy (ghazi):

hope that helps you

OpenStudy (ghazi):

theta is angle from horizontal = x

OpenStudy (anonymous):

mg sinx = ma sinx = a/g

OpenStudy (ghazi):

i think your answer will be = \[m*g* (\sqrt{a^2+g^2\cos^{2}x})* \cos x\]

OpenStudy (ghazi):

hmm..that's good ....sorry sometimes complex thinking never helps... :) good job

OpenStudy (anonymous):

thxxx

OpenStudy (ghazi):

:)